# Most Juntas Saturate the Hardcore Lemma

### v1

@@ Acknowledgements
We thank Guy Blanc and David Zuckerman for illuminating discussions. We also thank Jeffrey Champion, Sabee Grewal, Rocco Servedio, and anonymous reviewers for comments on a draft of this work.
@@ Appendix A Optimal Approximating Circuits: Proofs of Theorem 7 and Theorem 4
In this section, we give motivation and a proof of the fundamental result of Andreev, Clementi, and Rolim [ACR97]. The exposition here has nontrivial overlap with that of Trevisan [Tre09].
@@ A.1 Motivation
A classical probabilistic argument of Shannon says that there exists functions mapping $n$ bits to one bit that have circuit complexity $\Omega(2^{n}/n)$. What is remarkable is that this argument is tight; Lupanov’s theorem states that any function $f:\{0,1\}^{n}\to\{0,1\}$ can be computed in circuit size $O(2^{n}/n)$ [Lup71].
@@
Notice that Lemma 1 instantiated with Theorem 7 gives Theorem 4. Lemma 1 was proven in Section 3, so we now focus on the proof of Theorem 7. This was implicitly proven in [ACR97], but we give a shorter proof using Spielman codes [Spi96].
@@ A.2 Overview of Theorem 7
But are there asymptotically good codes encodable in linear circuit size? Yes: Spielman codes.
Theorem 8 ([Spi96]).

### v2

@@ Acknowledgements
We thank Guy Blanc, Geoffrey Mon and David Zuckerman for illuminating discussions. We also thank Jeffrey Champion, Sabee Grewal, Rocco Servedio, and anonymous reviewers for comments on a draft of this work.
@@ Appendix A Optimal Approximating Circuits: Proofs of Theorem 7 and Theorem 4
In this section, we give motivation and a proof of the fundamental result of Andreev, Clementi, and Rolim [1]. The exposition here has nontrivial overlap with that of Trevisan [20].
@@ A.1 Motivation
A classical probabilistic argument of Shannon says that there exists functions mapping $n$ bits to one bit that have circuit complexity $\Omega(2^{n}/n)$. What is remarkable is that this argument is tight; Lupanov’s theorem states that any function $f:\{0,1\}^{n}\to\{0,1\}$ can be computed in circuit size $O(2^{n}/n)$ [13].
@@
Notice that Lemma 1 instantiated with Theorem 7 gives Theorem 4. Lemma 1 was proven in Section 3, so we now focus on the proof of Theorem 7. This was implicitly proven in [1], but we give a shorter proof using asymptotically good codes encodable in linear time [7, 17].
@@ A.2 Overview of Theorem 7
But are there asymptotically good codes encodable in linear circuit size? Indeed, non-explicit constructions of such codes were known to exist since 1974, thanks to Gelfand, Dobrushin, and Pinsker [7]. Non-explicit constructions suffice for our application, but we mention that Spielman codes are explicit constructions of such a code [17].
Theorem 8 ([7, 17]).